# Graph-Informed Adversarial Modeling: Infimal Subadditivity of Interpolative Divergences

### 20 Mar 2026 v1

@@ 6.2.2 Ball-throwing trajectory
Figure 4: Fifty ball-throwing trajectory time series from various distributions. (a) Samples from the target distribution with $v_{0}\sim N(\mu_{v}=4,\sigma_{v}^{2}=3^{2})$. (b) Fifty generated trajectories from a graph-informed GAN with $14$ localized discriminators trained using an interpolative divergence with a Lipschitz constraint. (c) Fifty generated trajectories from a graph-agnostic GAN with a monolithic discriminator trained using an interpolative divergence with a Lipschitz constraint.
@@
Figure 5: Training diagnostics for the ball-trajectory experiment from a representative run; the graph-informed GAN with $14$ localized discriminators aligned to child–parent families in the Bayesian network is compared to the graph-agnostic GAN with a monolithic discriminator. (a) and (b) show the evolution of the KL variational loss equation 34 on the training and validation sets, for an interpolative (2020 runs in total) and classical divergence ($10$ runs in total), i.e., with and without Lipschitz constraints enforced via spectral normalization. (c) and (d) show the corresponding energy-distance diagnostics, both for the full trajectory distribution and for the recovered physics parameters $(\hat{\mathsf{g}},\hat{v}_{0})$, again for an interpolative and classical divergence.
@@ 6.3 Real-world Bayesian networks
We now turn to two real-world Bayesian networks in order to study graph-informed adversarial training beyond the synthetic continuous setting. The first example uses the Child network from Spiegelhalter (1992), a medium-sized Bayesian network consisting of $20$ categorical variables and $25$ directed edges ($230$ parameters) that describe clinical factors associated with birth asphyxia (Figure 6(a)). The second example uses the Earthquake network from Pearl (1988) (see also Korb and Nicholson (2010)), a smaller Bayesian network consisting of $5$ nodes and $4$ directed edges ($10$ parameters). In both cases, the network structure and CPTs are assumed to be known. These experiments are intended to test whether the graph-informed GAN continues to improve structural fidelity in discrete settings, and whether the resulting multi-discriminator training remains stable across repeated runs.
@@ 6.3.2 Child
The Child experiment assesses the graph-informed GAN on a medium-sized discrete Bayesian network in the theory-aligned child–parent setting and examines how the size of the localized discriminator families affects the trade-off between local and global fit of model predictions. We report results over $15$ independent runs.
@@
For Child, we write $M$ for the graph-agnostic GAN model with a monolithic discriminator, and $M_{a}$ for the graph-informed GAN model with localized discriminators acting on the families of ancestor depth $a$,
@@
The Child experiment shows that on a medium-sized discrete Bayesian network, graph-informed localization improves structural fidelity, as directly reflected in node-wise total variation, local AUC, and log-likelihood. Broader neighborhoods partly recover global fit, but the main advantage of the graph-informed GAN in this setting is better alignment with the conditional structure encoded by the Bayesian network.
@@
Figure 7: Training and validation variational loss (interpolative JS-Wasserstein) for the real-world discrete Bayesian network experiments. (a) Compares the graph-agnostic GAN model $M$ with the graph-informed GAN models $M_{1}$ (child–parent families) and $M_{3}$ (ancestor depth-33 families) on Child over $15$ runs. (b) Compares the graph-agnostic GAN and the graph-informed GAN on Earthquake over $85$ runs.
@@
Figure 8: Validation diagnostics for the real-world discrete Bayesian network experiments (single run highlighted). The Child experiment compares $M$, $M_{1}$, and $M_{3}$, corresponding to graph-agnostic GAN and two graph-informed GANs with child–parent and ancestor depth-3 families over $15$ runs. The Earthquake experiment compares graph-agnostic and graph-informed GANs over $85$ runs. (a) and (b) show global validation metrics, namely energy distance and average log-likelihood. (c) and (d) show representative discriminator AUC values.
@@ 6.3.3 Earthquake
Figure 7(b) again shows smoother training dynamics for the graph-informed GAN than for the graph-agnostic GAN. The difference is less dramatic than for Child, but the graph-informed model still exhibits less erratic variation in the training and validation objectives. In the global validation metrics shown in Figure 8(b), the two models are broadly comparable in terms of their central tendency. This is consistent with the fact that the full joint distribution is small enough that the monolithic discriminator is not severely disadvantaged.
@@
This provides a useful boundary case for the broader conclusions of the paper. When the joint space is small, the graph-informed GAN still offers a practical advantage by making training more stable and reducing the frequency of poor outcomes across repeated runs.
@@ 7 Discussion
We have introduced an infimal subadditivity principle for interpolative divergences on Bayesian networks and used it to give a variational justification for graph-informed adversarial learning. In particular, Theorems 5.4, 5.6, and 5.7 show that, when the target distribution factorizes according to a known DAG, a global discrepancy objective functional can be controlled by an average of family-level objective functionals aligned with the graph. This provides a principled basis for replacing a graph-agnostic GAN with a monolithic discriminator by a graph-informed GAN with localized discriminators, without requiring the optimizer itself to factorize according to the same graph. The same variational framework applies to $(f,\Gamma)$-divergences, $\Gamma$-IPMs, and proximal optimal transport divergences, and applies to natural discriminator classes including bounded measurable and Lipschitz-bounded functions.
@@ Appendix B Supporting material
$\begin{split}&W_{1}\left(Q(x_{3}\mid X_{2}=x_{2}),Q(dx_{3}\mid X_{2}=x^{\prime}_{2})\right)\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{X}_{3}}f(x_{3})Q(dx_{3}\mid X_{2}=x_{2})-\int_{\mathcal{X}_{3}}f(x_{3})Q(dx_{3}\mid X_{2}=x^{\prime}_{2})\right\}\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{X}_{3}}f(x_{3})g(x_{2},\cdot)_{\#}\,\nu(dz)-\int_{\mathcal{X}_{3}}f(x_{3})g(x^{\prime}_{2},\cdot)_{\#}\,\nu(dz)\right\}\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{Z}}f(g(x_{2},z))\,\nu(dz)-\int_{\mathcal{Z}}f(g(x^{\prime}_{2},z))\,\nu(dz)\right\}\\ &\qquad\leq\int_{\mathcal{Z}}\|g(x_{2},z)-g(x^{\prime}_{2},z)\|\,d\nu(z)\\ &\qquad\leq Cd_{\mathcal{X}_{2}}(x_{2},x^{\prime}_{2})\end{split}$
The proof of $W_{1}\left(Q(dx_{2}\mid X_{1}=x_{1}),Q(dx_{2}\mid X_{1}=x^{\prime}_{1})\right)\leq Cd_{\mathcal{X}_{1}}(x_{1},x^{\prime}_{1})$ goes similarly.
We now formalize the preceding observation in a proposition that is used to verify condition (C) for Lipschitz-bounded discriminator classes.
@@
Let $Q\in\mathcal{P}^{G}$ and $G$ is given by equation 17. Suppose that the map $\bm{x}_{\mathrm{Pa}(i)}\to Q(d\bm{x}_{V(i)}\mid\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)})$, $i\in V$ is $C$-Lipschitz with respect to $1$-Wasserstein metric, respectively, i.e., for any $\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)}\in\mathcal{X}_{\mathrm{Pa}(i)}$
$W_{1}(Q(dx_{i}\mid\bm{X}_{\mathrm{Pa}(i)}=\bm{x}_{\mathrm{Pa}(i)}),Q(dx_{i}\mid\bm{X}_{\mathrm{Pa}(i)}=\bm{x}^{\prime}_{\mathrm{Pa}(i)}))\leq C\cdot d_{\mathcal{X}_{\mathrm{Pa}(i)}}(\bm{x}_{\mathrm{Pa}(i)},\bm{x}^{\prime}_{\mathrm{Pa}(i)})\,.$ (39)
If $\Gamma^{L}={\rm{Lip}}_{b}^{L}(\mathcal{X})$, then condition (C) holds with $L_{i}=L+C$.
@@
Let $\gamma\in\Gamma$. To show that (C) holds, we prove that for any pairs $(x_{i},\bm{x}_{\mathrm{Pa}(i)})$ and $(x^{\prime}_{i},\bm{x}^{\prime}_{\mathrm{Pa}(i)})\in\mathcal{X}_{i\,\cup\,\mathrm{Pa}(i)}$
$|I_{i,\mathrm{Pa}(i)}[\gamma](x_{i},\bm{x}_{\mathrm{Pa}(i)})-I_{i,\mathrm{Pa}(i)}[\gamma](x^{\prime}_{i},\bm{x}^{\prime}_{\mathrm{Pa}(i)})|\leq L\cdot d_{\mathcal{X}_{i}\times\mathcal{X}_{\mathrm{Pa}(i)}}\left((x_{i},\bm{x}_{\mathrm{Pa}(i)}),(x^{\prime}_{i},\bm{x}^{\prime}_{\mathrm{Pa}(i)})\right).$
We start from the left-hand side of the above inequality for $i=3$ and $\mathrm{Pa}(i)=\mathrm{Pa}(3)=2$, and bound as follows: Let $\gamma(\bm{x})=\gamma(x_{1},x_{2},x_{3})$
$|I_{3,2}[\gamma](x_{3},x_{2})-I_{3,2}[\gamma](x^{\prime}_{3},x^{\prime}_{2})|$
$\hskip 71.13188pt=\Big|\int_{\mathcal{X}_{1}}\gamma(\bm{x})Q(dx_{1}\mid X_{2}=\bm{x}_{2})-\int_{X_{1}}\gamma(\bm{x}^{\prime})Q(dx_{1}\mid X_{2}=\bm{x}^{\prime}_{2})\Big|$
$\hskip 71.13188pt=\int_{\mathcal{X}_{1}}\int_{\mathcal{X}_{1}}|\gamma(\bm{x})-\gamma(\bm{x}^{\prime})|\,Q(dx_{1}\mid X_{2}=x_{2})Q(dx_{1}|X_{2}=x^{\prime}_{2})$
$\hskip 71.13188pt=\int_{\mathcal{X}_{1}}\int_{\mathcal{X}_{1}}Ld_{\mathcal{X}}(\bm{x},\bm{x}^{\prime})\pi_{x_{2},x^{\prime}_{2}}(dx_{1})$
$\hskip 71.13188pt\leq L\cdot d_{\mathcal{X}_{3}\times\mathcal{X}_{2}}\left((x_{2},x_{3}),(x^{\prime}_{2},x^{\prime}_{3})\right)$
$\hskip 113.81102pt+W_{1}(Q(dx_{1}\mid X_{2}=x_{2}),Q(dx_{1}\mid X_{2}=x^{\prime}_{2}))$
where $\pi_{x_{2},x_{2}^{\prime}}$ is any coupling of $Q(dx_{1}\mid X_{2}=x_{2})$ and $Q(dx_{1}\mid X_{2}=x^{\prime}_{2})$. The result follows immediately. We work similarly $I_{1}[\gamma]$ and $I_{2,1}[\gamma]$. ∎

### 02 Apr 2026 v2

@@ 6.2.2 Ball-throwing trajectory
Figure 4: Fifty ball-throwing trajectory time series from various distributions. (a) Samples from the target distribution with $v_{0}\sim N(\mu_{v}=4,\sigma_{v}^{2}=3^{2})$. (b) Fifty generated trajectories from a GiGAN with $14$ localized discriminators trained using an interpolative divergence with a Lipschitz constraint. (c) Fifty generated trajectories from a graph-agnostic GAN with a monolithic discriminator trained using an interpolative divergence with a Lipschitz constraint.
@@
Figure 5: Training diagnostics for the ball-trajectory experiment from a representative run; the GiGAN with $14$ localized discriminators aligned to child–parent families in the Bayesian network is compared to the graph-agnostic GAN with a monolithic discriminator. (a) and (b) show the evolution of the KL variational loss equation 37 on the training and validation sets, for an interpolative (2020 runs in total) and classical divergence ($10$ runs in total), i.e., with and without Lipschitz constraints enforced via spectral normalization. (c) and (d) show the corresponding energy-distance diagnostics, both for the full trajectory distribution and for the recovered physics parameters $(\hat{\mathsf{g}},\hat{v}_{0})$, again for an interpolative and classical divergence.
@@ 6.3 Real-world Bayesian networks
We now turn to two real-world Bayesian networks in order to study graph-informed adversarial training beyond the synthetic continuous setting. The first example uses the Child network from Spiegelhalter (1992), a medium-sized Bayesian network consisting of $20$ categorical variables and $25$ directed edges ($230$ parameters) that describe clinical factors associated with birth asphyxia (Figure 6(a)). The second example uses the Earthquake network from Pearl (1988) (see also Korb and Nicholson (2010)), a smaller Bayesian network consisting of $5$ nodes and $4$ directed edges ($10$ parameters). In both cases, the network structure and CPTs are assumed to be known. These experiments are intended to test whether the GiGAN continues to improve structural fidelity in discrete settings, and whether the resulting multi-discriminator training remains stable across repeated runs.
@@ 6.3.2 Child
The Child experiment assesses the GiGAN on a medium-sized discrete Bayesian network in the theory-aligned child–parent setting and examines how the size of the localized discriminator families affects the trade-off between local and global fit of model predictions. We report results over $15$ independent runs.
@@
For Child, we write $M$ for the graph-agnostic GAN model with a monolithic discriminator, and $M_{a}$ for the GiGAN model with localized discriminators acting on the families of ancestor depth $a$,
@@
The Child experiment shows that on a medium-sized discrete Bayesian network, graph-informed localization improves structural fidelity, as directly reflected in node-wise total variation, local AUC, and log-likelihood. Broader neighborhoods partly recover global fit, but the main advantage of the GiGAN in this setting is better alignment with the conditional structure encoded by the Bayesian network.
@@
Figure 7: Training and validation variational loss (interpolative JS-Wasserstein) for the real-world discrete Bayesian network experiments. (a) Compares the graph-agnostic GAN model $M$ with the GiGAN models $M_{1}$ (child–parent families) and $M_{3}$ (ancestor depth-33 families) on Child over $15$ runs. (b) Compares the graph-agnostic GAN and the GiGAN on Earthquake over $85$ runs.
@@
Figure 8: Validation diagnostics for the real-world discrete Bayesian network experiments (single run highlighted). The Child experiment compares $M$, $M_{1}$, and $M_{3}$, corresponding to graph-agnostic GAN and two GiGANs with child–parent and ancestor depth-3 families over $15$ runs. The Earthquake experiment compares graph-agnostic and GiGANs over $85$ runs. (a) and (b) show global validation metrics, namely energy distance and average log-likelihood. (c) and (d) show representative discriminator AUC values.
@@ 6.3.3 Earthquake
Figure 7(b) again shows smoother training dynamics for the GiGAN than for the graph-agnostic GAN. The difference is less dramatic than for Child, but the graph-informed model still exhibits less erratic variation in the training and validation objectives. In the global validation metrics shown in Figure 8(b), the two models are broadly comparable in terms of their central tendency. This is consistent with the fact that the full joint distribution is small enough that the monolithic discriminator is not severely disadvantaged.
@@
This provides a useful boundary case for the broader conclusions of the paper. When the joint space is small, the GiGAN still offers a practical advantage by making training more stable and reducing the frequency of poor outcomes across repeated runs.
@@ 7 Discussion
We have introduced an infimal subadditivity principle for interpolative divergences on Bayesian networks and used it to give a variational justification for graph-informed adversarial learning. In particular, Theorems 5.3, 5.5, and 5.6 show that, when the target distribution factorizes according to a known DAG, a global discrepancy objective functional can be controlled by an average of family-level objective functionals aligned with the graph. This provides a principled basis for replacing a graph-agnostic GAN with a monolithic discriminator by a GiGAN, a graph-informed GAN with localized discriminators trained under an interpolative divergence objective, without requiring the optimizer itself to factorize according to the same graph. The same variational framework applies to $(f,\Gamma)$-divergences, $\Gamma$-IPMs, and proximal optimal transport divergences, and applies to natural discriminator classes including bounded measurable and Lipschitz-bounded functions.
@@ Appendix B Supporting material
$\begin{split}&W_{1}\left(Q(dx_{3}|x_{2}),Q(dx_{3}|x^{\prime}_{2})\right)\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{X}_{3}}f(x_{3})Q(dx_{3}|x_{2})-\int_{\mathcal{X}_{3}}f(x_{3})Q(dx_{3}|x^{\prime}_{2})\right\}\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{X}_{3}}f(x_{3})(g(x_{2},\cdot)_{\#})\,\nu(dz)-\int_{\mathcal{X}_{3}}f(x_{3})(g(x^{\prime}_{2},\cdot)_{\#})\,\nu(dz)\right\}\\ &\qquad=\sup_{f\in\mathrm{Lip}_{b}^{1}(\mathcal{X}_{3})}\left\{\int_{\mathcal{Z}}f(g(x_{2},z))\,\nu(dz)-\int_{\mathcal{Z}}f(g(x^{\prime}_{2},z))\,\nu(dz)\right\}\\ &\qquad\leq\int_{\mathcal{Z}}\|g(x_{2},z)-g(x^{\prime}_{2},z)\|\,d\nu(z)\\ &\qquad\leq Cd_{\mathcal{X}_{2}}(x_{2},x^{\prime}_{2})\end{split}$
The proof of $W_{1}\left(Q(dx_{2}|x_{1}),Q(dx_{2}|x^{\prime}_{1})\right)\leq Cd_{\mathcal{X}_{1}}(x_{1},x^{\prime}_{1})$ goes similarly.
We now formalize the preceding observation in a proposition that is used to verify condition (C) for Lipschitz-bounded discriminator classes. For the next proposition, we denote
$K_{i}(d\bm{x}_{[i\,\cup\,\mathrm{Pa}(i)]^{c}}|\bm{x}_{i\,\cup\,\mathrm{Pa}(i)}):=Q_{V^{(i-1)}|\mathrm{Pa}(i)}(d\bm{x}_{V^{(i-1)}})\prod_{j=i+1}^{n}Q_{j|\mathrm{Pa}(j)}(dx_{j}),$
where we remind that $V^{(i)}=\{1,\dots,i\}\setminus\mathrm{Pa}(i+1)$. The next proposition is stated for DAG equation 18 for simplicity, it can be easily generalized to an arbitrary DAG.
@@
Let $G$ be the chain $X_{1}\to X_{2}\to X_{3}$. Assume that
$W_{1}(Q_{2|1}(\cdot|x_{1}),Q_{2|1}(\cdot|x_{1}^{\prime}))\leq C_{2}\,d_{\mathcal{X}_{1}}(x_{1},x_{1}^{\prime}),$
$W_{1}(Q_{3|2}(\cdot|x_{2}),Q_{3|2}(\cdot|x_{2}^{\prime}))\leq C_{3}\,d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime}).$
Then $K_{i}$ is Lipschitz with Lipschitz constant $C$, with respect to the $1$-Wasserstein distance.
@@
We prove the result for $i=1$ (the proof for $i=2,3$ follows the same steps), that is, we prove that
$W_{1}(K_{1}(\cdot|x_{1}),K_{1}(\cdot|x_{1}^{\prime}))\leq(1+C_{3})C_{2}\,d_{\mathcal{X}_{1}}(x_{1},x_{1}^{\prime}).$
Fix $x_{1},x_{1}^{\prime}$. Let $\pi_{22^{\prime}}$ be an optimal coupling of $Q_{2|1}(\cdot|x_{1})$ and $Q_{2|1}(\cdot|x_{1}^{\prime})$, so that
$\int d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime})\,\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime})=W_{1}((Q_{2|1}(\cdot|x_{1}),Q_{2|1}(\cdot|x_{1}^{\prime})).$
For each $(x_{2},x_{2}^{\prime})$, let $\pi_{33^{\prime}}^{x_{2},x_{2}^{\prime}}$ be an optimal coupling of $Q_{3|2}(\cdot|x_{2})$ and $Q_{3|2}(\cdot|x_{2}^{\prime})$, i.e.,
$\int d_{\mathcal{X}_{3}}(x_{3},x_{3}^{\prime})\,\pi_{33^{\prime}}^{x_{2},x_{2}^{\prime}}(dx_{3},dx_{3}^{\prime})=W_{1}(Q_{3|2}(\cdot|x_{2}),Q_{3|2}(\cdot|x_{2}^{\prime})).$
Define a coupling $\Pi$ of $K_{1}(\cdot|x_{1})$ and $K_{1}(\cdot|x_{1}^{\prime})$ by
$\Pi(dx_{2},dx_{3},dx_{2}^{\prime},dx_{3}^{\prime})=\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime})\,\pi_{33^{\prime}}^{x_{2},x_{2}^{\prime}}(dx_{3},dx_{3}^{\prime}).$
It is straightforward to verify that $\Pi$ has the correct marginals.
By definition of $W_{1}$ and using the product metric,
$W_{1}(K_{1}(\cdot|x_{1}),K_{1}(\cdot|x_{1}^{\prime}))$ $\leq$ $\int\bigl(d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime})+d_{\mathcal{X}_{3}}(x_{3},x_{3}^{\prime})\bigr)\,d\Pi$
$=$ $\int d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime})\,\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime})$
$\quad+\int\int\left(\int d_{\mathcal{X}_{3}}(x_{3},x_{3}^{\prime})\,\pi_{33^{\prime}}^{x_{2},x_{2}^{\prime}}(dx_{3},dx_{3}^{\prime})\right)\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime}).$
Using optimality of the couplings, this yields
$W_{1}(K_{1}(\cdot|x_{1}),K_{1}(\cdot|x_{1}^{\prime}))$ $\leq$ $W_{1}(Q_{2|1}(\cdot|x_{1}),Q_{2|1}(\cdot|x_{1}^{\prime}))$
$\quad+\int W_{1}(Q_{3|2}(\cdot|x_{2}),Q_{3|2}(\cdot|x_{2}^{\prime}))\,\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime}).$
By the Lipschitz assumption on $Q_{3|2}$,
$W_{1}(Q_{3|2}(\cdot|x_{2}),Q_{3|2}(\cdot|x_{2}^{\prime}))\leq C_{3}\,d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime}),$
so
$\int W_{1}(Q_{3|2}(\cdot|x_{2}),Q_{3|2}(\cdot|x_{2}^{\prime}))\,\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime})\leq C_{3}\int d_{\mathcal{X}_{2}}(x_{2},x_{2}^{\prime})\,\pi_{22^{\prime}}(dx_{2},dx_{2}^{\prime}).$
Therefore,
$W_{1}(K_{1}(\cdot|x_{1}),K_{1}(\cdot|x_{1}^{\prime}))$
$\leq(1+C_{3})W_{1}(Q_{2|1}(\cdot|x_{1}),Q_{2|1}(\cdot|x_{1}^{\prime}))\leq(1+C_{3})C_{2}\,d_{\mathcal{X}_{1}}(x_{1},x_{1}^{\prime}),$
which proves the claim. ∎
Proposition B.2.
Let $Q\in\mathcal{P}^{G}$ and let $G$ be an arbitrary DAG. Assume that, for each $i\in V$, the kernel $K_{i}(\cdot|\bm{x}_{i\cup\mathrm{Pa}(i)})$ is $C$-Lipschitz with respect to $W_{1}$, i.e.,
$W_{1}(K_{i}(\cdot|\bm{x}_{i\cup\mathrm{Pa}(i)}),K_{i}(\cdot|\bm{x}^{\prime}_{i\cup\mathrm{Pa}(i)}))\leq C\,d_{\mathcal{X}_{i\cup\mathrm{Pa}(i)}}(\bm{x}_{i\cup\mathrm{Pa}(i)},\bm{x}^{\prime}_{i\cup\mathrm{Pa}(i)}).$ (42)
If $\Gamma^{L}=\mathrm{Lip}_{b}^{L}(\mathcal{X})$, then condition (C) holds with
$L_{i}=L(1+C)>L,\qquad i\in V.$
Proof.
Fix $i\in V$ and let
$u=\bm{x}_{i\cup\mathrm{Pa}(i)},\qquad u^{\prime}=\bm{x}^{\prime}_{i\cup\mathrm{Pa}(i)}.$
For $\gamma\in\Gamma^{L}$, define
$I_{i,\mathrm{Pa}(i)}[\gamma](u)=\int_{\mathcal{X}_{[i\cup\mathrm{Pa}(i)]^{c}}}\gamma(u,z)\,K_{i}(dz|u).$
Let $\Pi$ be any coupling of $K_{i}(\cdot|u)$ and $K_{i}(\cdot|u^{\prime})$. Then
$|I_{i,\mathrm{Pa}(i)}[\gamma](u)-I_{i,\mathrm{Pa}(i)}[\gamma](u^{\prime})|$ $=$ $\left|\int\gamma(u,z)\,K_{i}(dz|u)-\int\gamma(u^{\prime},z^{\prime})\,K_{i}(dz^{\prime}|u^{\prime})\right|$
$=$ $\left|\int(\gamma(u,z)-\gamma(u^{\prime},z^{\prime}))\,\Pi(dz,dz^{\prime})\right|$
$\leq$ $\int|\gamma(u,z)-\gamma(u^{\prime},z^{\prime})|\,\Pi(dz,dz^{\prime}).$
Since $\gamma$ is $L$-Lipschitz,
$|\gamma(u,z)-\gamma(u^{\prime},z^{\prime})|\leq L\Big(d_{\mathcal{X}_{i\cup\mathrm{Pa}(i)}}(u,u^{\prime})+d_{\mathcal{X}_{[i\cup\mathrm{Pa}(i)]^{c}}}(z,z^{\prime})\Big).$
Therefore,
$|I_{i,\mathrm{Pa}(i)}[\gamma](u)-I_{i,\mathrm{Pa}(i)}[\gamma](u^{\prime})|$ $\leq L\,d_{\mathcal{X}_{i\cup\mathrm{Pa}(i)}}(u,u^{\prime})$
$\quad+L\int d_{\mathcal{X}_{[i\cup\mathrm{Pa}(i)]^{c}}}(z,z^{\prime})\,\Pi(dz,dz^{\prime}).$
Choosing $\Pi$ optimally and using equation 42, we obtain
$|I_{i,\mathrm{Pa}(i)}[\gamma](u)-I_{i,\mathrm{Pa}(i)}[\gamma](u^{\prime})|$ $\leq L\,d_{\mathcal{X}_{i\cup\mathrm{Pa}(i)}}(u,u^{\prime})+L\,W_{1}\!\left(K_{i}(\cdot|u),K_{i}(\cdot|u^{\prime})\right)$
$\leq L(1+C)\,d_{\mathcal{X}_{i\cup\mathrm{Pa}(i)}}(u,u^{\prime}).$
Boundedness is straightforward. Hence
$I_{i,\mathrm{Pa}(i)}[\gamma]\in\mathrm{Lip}_{b}^{L(1+C)}(\mathcal{X}_{i\cup\mathrm{Pa}(i)}),$
which proves the desired result. ∎